Please help me to solve these equations using the elimination method?
0.06x+0.05y=0.07
0.4x-0.3y=1.1

-4.5x+7.5y=6
-x+1.5y=5

17x+21y=1
17x-5y=6
Thank you

Answer by scott8148(6628)

0.06x+0.05y=0.07
0.4x-0.3y=1.1
__ multiply 1st equation by 6 and add to 2nd equation

-4.5x+7.5y=6
-x+1.5y=5
__ multiply 2nd equation by -5 and add to 1st equation

17x+21y=1
17x-5y=6
__ subtract 2nd equation from 1st equation


after you have solved for one variable, substitute the value into one of the equations to find the other variable